# Optimizing Probabilistic Conformal Prediction with Vectorized Non-Conformity Scores

### v1

@@ 3.2 Efficiency-Enhanced Prediction Set
Efficient Algorithm for Quantile Optimization. Solving problem (8) becomes computationally infeasible for larger values of $K$ (e.g., $K\geq 5$) due to the exponential growth of the search space. To tackle this challenge, we introduce an iterative algorithm that approximates the optimal quantile vector in linear time by simplifying the search process. The key idea is to activate one quantile at a time while setting others to default values, and then perform pairwise trade-offs between quantiles of different ranks to enhance efficiency. Specifically, we initialize the algorithm by setting $\beta_{r}=0$ for a selected rank $r$, with all other $\beta_{s}=1$ for $s\neq r$, effectively reducing the non-conformity score vector to a scalar case. We incrementally increase $\beta_{r}$ by a small step size $\epsilon$ until the empirical coverage just exceeds the target level $1-\alpha$. Subsequently, we iteratively adjust $\beta_{r}$ and other $\beta_{s}$ by trading off small increments and decrements while maintaining the coverage constraint, accepting adjustments that improve efficiency. This process has a computational complexity of $\mathcal{O}(Kn)$ (or $\mathcal{O}(K^{2}n)$ when iterating over all positions) and empirically yields an approximate solution $\hat{\beta}$ that achieves superior efficiency compared to baseline methods. The details of the algorithm can be found in Appendix C.
@@ 4 Experiments
In this section, we evaluate the empirical performance of PCP-VCR compared with several state-of-the-art baseline methods on five synthetic data, one semi-synthetic data, and four real-world datasets. We demonstrate the superior efficiency and coverage validity of PCP-VCR, especially in handling complex, multimodal, and heterogeneous distributions.
@@ 4.1 Synthetic Data
We present the summary of the synthetic data results averaged over 100 experiments in Table 1 and highlight the most efficient method for each data. Overall, all methods achieve valid coverage rates, while PCP-VCR produces the most efficient prediction set. The visualizations of the prediction sets of different methods on these data are presented in Appendix E. We can observe that PCP and PCP-VCR can consistently provide discontinuous prediction sets and, thus are much more efficient than other methods for multimodal settings. Compared with PCP, PCP-VCR provides a more flexible and locally-adaptive prediction set, i.e., the coverage region of PCP-VCR is larger in the high-density region and smaller in the low-density region, especially in the Mix-Gaussian setting. This is the benefit that vectorized non-conformity scores provide compared with a scalar.
Figure 2: Comparison of MNIST images sampled from the prediction set provided by PCP-VCR and PCP, respectively. Each row corresponds to different values of $X$. We observe that samples of PCP-VCR align more closely with the target $X$ compared to PCP. It indicates that PCP-VCR provides more efficient prediction sets with fewer wrong images covered.
4.2 MNIST: Digit Image Generation
We design a semi-synthetic experiment with the MNIST image dataset (Deng,, 2012) to assess the applicability and interpretability of PCP-VCR to high-dimensional data. Conformal prediction in high-dimensional data, such as images, texts, and audio, presents a unique challenge for interpreting prediction sets, as it is inapplicable to iterate over the entire space of target variables.
Consider the setting that given $X\in\{0,\dots,9\}$, the corresponding Y is an image of digit $X$. We first train a variational-autoencoder (VAE) model and create the prediction set in the embedding space. For the conditional sampling process, we assume we have access to an independent sampler. For any given $X$, it will output random images of digit $X$ that are not presented in the training/calibration/testing process to ensure independence. We then embed the sampled images with the trained encoder and construct the prediction set in the embedding space. Ahead of training, we split part of the data for the sampler so it is independent of the following process. The reason that we design this sampler instead of training a conditional VAE (cVAE) is that the embedding space of VAE tends to be more clustered than cVAE, thus the prediction set can be better demonstrated for illustrative purposes.
Figure 2 shows samples from the prediction set provided by PCP and PCP-VCR with $K=40$ conditioning on different $X$ values in each row. For each $X$, we randomly sample 9 images from the prediction set of PCP and PCP-VCR. Both of the methods achieve valid coverage rates. We can observe that the sampled images of PCP-VCR align closer with the target image for each $X$, indicating a tighter and more efficient prediction set of PCP-VCR. Quantitatively, as it is not applicable to directly measure the size of the prediction set, we propose to evaluate the miscoverage rate. It measures the empirical probability of an incorrect image falling into the prediction set. Under $K=40$, PCP-VCR attains an average miscoverage rate of $0.16\pm 0.02$ (standard error) over 30 experiments, compared to $0.22\pm 0.03$ for PCP. In Figure 3, we present the comparison of the miscoverage rate of PCP and PCP-VCR under different choices of $K$. It demonstrates that PCP-VCR has a lower miscoverage rate than PCP consistently. This also explains why the random images of PCP-VCR aligns more closely with the target as shown in Figure 2
Figure 3: Comparison of miscoverage rate of PCP and PCP-VCR for MNIST data under different $K$. The blue and red bars represent PCP and PCP-VCR results, respectively, with black lines showing the standard error over 30 experiments.
@@ 4.2 MNIST: Digit Image Generation
  • 1
    Each cell contains the averaged metric value along with the standard error over 30 independent experiments.
(a) PCP-VCR
@@ 4.3 Real Data
We apply PCP-VCR along with baseline methods on four real-world datasets: blog feedback (Blog) Buza, (2014), facebook comment volumne Singh, (2015) variant 1 (Fb1) and variant 2 (Fb2), and New York City bike sharing (Bike). For each dataset, we randomly sample $30,000$ samples for training, $1,000$ samples for calibration and $1,000$ for testing. We set $K=50$ for the Bike dataset, and $K=20$ for the rest datasets. The summary of the real data results averaging over 30 experiments is presented in Table 2. Overall, PCP-VCR still performs the best, and it achieves the best efficiency while maintaining the coverage rate. It is noted that for the first three datasets, R2CCP suffers from the undercover issue and has a large coverage rate variance across experiments.
We present a visual demonstration of the prediction set generated by our method, PCP, CHR and R2CCP in Figure 4, respectively. The blue region indicates the prediction set of each method and the red star marker indicates the target location. Both PCP and our method deliver discontinuous prediction regions, whereas CHR and R2CCP provide rectangle-shaped continuous sest. We also notice that, compared with PCP, PCP-VCR assigns larger circles’ radius to high-density regions while nearly vanishing the circles’ radius in low-density regions. Specifically, the circle at the top of the map lies in a low-density region, it almost shrinks to a point for PCP-VCR, whereas it remains large for PCP. Thus, PCP-VCR leads to significant improvements in flexibility and efficiency of prediction set.
@@ 5 Discussion and Conclusion
In this paper, we proposed PCP-VCR, a novel conformal prediction method for conditional generative models utilizing vectorized non-conformity scores with ranked samples to enhance the efficiency of the prediction set. By ranking the conditional samples, our method explores and captures the structure among the random samples and incorporates it into non-conformity score vector. By separately optimizing each entry of the non-conformity score vector, it enables the creation of a flexible and data-adaptive prediction set. Specifically, PCP-VCR outputs discontinuous prediction set in the form of a union of regions centered at random samples, where the radius of each region is adapted to the model’s uncertainty level of each sample. Additionally, we proposed a computationally efficient algorithm for approximating the optimal quantile vector, ensuring the method practical even for large datasets.
Through extensive experiments on both synthetic and real-world datasets, PCP-VCR consistently demonstrated superior coverage efficiency while maintaining coverage validity. In the MNIST experiment, we showcased the potential power of PCP-VCR for conformal prediction in general generative model settings, illustrating its ability to improve the reliability of predictions in modern high-stakes applications.
@@ Appendix A Average $m$-Nearest Neighbor Distance
Appendix B Proof of Theorem 3.1
@@
We can rewrite the constraint in Equation 8 as
$\begin{split}\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)&=\frac{1}{n+1}\sum_{i=1}^{n}\max_{r=1,\dots,K}\mathbbm{1}\left\{E_{i,r}\leq Q_{r}(\beta_{r})\right\}\\ &=\frac{1}{n+1}\sum_{i=1}^{n}(1-\mathbbm{1}\{E_{i}\succ Q(\beta)\})\\ &\geq 1-\alpha,\end{split}$ (10)
where we let $Q(\beta)=(Q_{1}(\beta_{1}),\dots,Q_{K}(\beta_{K}))^{T}$ denote the vector of quantiles for simplicity. We can then observe the following inequality based on the last inequality in Equation 10,
$\begin{split}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{E_{i}\succ Q(\beta)\}<\alpha\end{split}$ (11)
We can then define the empirical multivariate quantile function as
$\hat{Q}(\{E_{1},\dots,E_{n}\},1-\alpha)$ $=\operatorname*{argmin}_{\beta\in\mathbb{R}^{K}}\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}$ (12)
s.t. $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{E_{i}\succ Q(\beta)\}<\alpha$ (13)
We let $\beta^{*}=\hat{Q}(\{E_{1},\dots,E_{n}\},1-\alpha)$, then for a testing data pair $(X_{n+1},Y_{n+1})$ that is exchangeable with $\mathcal{D}$, we have $\mathbb{P}(E_{n+1}\succ Q(\beta^{*}))<\alpha,$ by definition of $Q(\beta^{*})$. We observe that its complement set achieves the target coverage, i.e,
@@
$\cup_{r=1}^{K}\{E_{n+1,r}\leq Q_{r}(\beta^{*}_{r})\}\Longleftrightarrow\{Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*})\}.$
@@ Appendix C Efficient Algorithm for Quantile Optimization
The key idea of the approximation method is motivated by the “trade-off” between quantiles for different ranks. Intuitively, if one quantile is decreased then another quantile has to be increased to maintain coverage validity. During this trade-off, the efficiency changes accordingly. We present a simple demonstration of this idea for $K=2$ in Figure 5. The colored region represents the feasible region in 8 with darker color indicating higher efficiency, and the red point shows the position of $\beta^{*}$. We can observe ($i$) a step-wise pattern on the boundary of the feasible region and ($ii$) $\beta^{*}$ appears on the boundary. The step-wise pattern corresponds to the trade-off between $\beta_{1}$ and $\beta_{2}$, which leads to different efficiency. Additionally, the feasible region is in a convex shape, also indicating the trade-off relationship between quantiles. Thus to approximate the optimal solution, we only need to find the boundary of the feasible region, and iteratively do the trade-offs between different quantiles.
Figure 5: Feasible region and optimal solution of optimization problem in Equation 8 for $K=2$. The colored region shows the feasible region of $\beta$ with darker color indicating higher efficiency and the red point shows the position of $\beta^{*}$.
To be more specific, the algorithm initializes a $\beta_{k}=0$ and sets all the other entries of $\beta$ to be 1. Then, we gradually increase $\beta_{k}$ until the empirical coverage is just achieved above the target coverage. The intuition behind such initialization is that it reduces the non-conformity score vector to the simple scalar case, as observed in the following equation,
@@
for $\beta_{k}\neq 1$ and $\beta_{j}=1$ for all $j\neq k$. Thus there exists such $\beta_{k}$ that $\beta$ can still satisfy empirical coverage with only $\beta_{k}\neq 1$. Let $\epsilon>0$ denote a small increment, we then randomly pick another entry $\beta_{j}$ for $j\neq k$ and do the trade-off between $\beta_{j}$ and $\beta_{k}$. We set $\beta_{k}=\beta_{k}+\epsilon$, and decrease $\beta_{j}$ until the empirical coverage is satisfied. Elementally, this process mimics the step-wise pattern, where we increase $\beta_{k}$ and decrease $\beta_{j}$ to do the trade-off and examine if the efficiency is improved. We iterate this process for different $j$ until reach the budget. We show that empirically, the approximated solution $\hat{\beta}$ still achieves superior efficiency than baseline methods. We refer to Algorithm 2 for the complete details of the empirical approximation method. As a remark for practice, we set $\epsilon=1/n$ as default. But in cases, if the quantiles are not updated after several iterations, one may consider increasing $\epsilon$ for a larger trade-off.
@@
3:  while $\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)>1-\alpha$ do
4:     $\beta_{k}\leftarrow\beta_{k}+\epsilon$;
@@
11:     while $\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)<1-\alpha$ do
@@
19:        Set $k$-th and $j$-th entry of $\beta$ to $\beta_{k}$ and $\beta_{j}$;
@@
Figure 7: Comparison of prediction set ($\alpha=0.1$) generated by different methods on $2$-dimensional synthetic mixture Gaussian data. The mixture Gaussian data is generated with 2 components with weights 0.7 and 0.3 from left to right, respectively. The blue shade indicates the true conditional density and the rectangles and circles denote the prediction sets of different methods. In the left subplot, we present the rectangle-shaped prediction set generated by CQR, CHR and R2CCP. In the right subplot, we present the circle-shaped prediction regions provided by PCP-VCR and PCP.
Figure 8: Ratio of PCP prediction set length over PCP-VCR prediction set length under different $K=3,5,10,15,20$ and number of calibration data for S-shape data. The blue curve represents the mean, and the shaded area indicates the standard error of the ratio. We observe that the ratio remains consistently above 1, demonstrating that PCP-VCR consistently achieves better efficiency.

### v2

@@ 3.2 Efficiency-Enhanced Prediction Set
Efficient Algorithm for Quantile Optimization. Solving problem (6) becomes computationally infeasible for larger values of $K$ (e.g., $K\geq 5$) due to the exponential growth of the search space. To tackle this challenge, we introduce an iterative algorithm that approximates the optimal quantile vector in linear time by simplifying the search process. The key idea is to activate one quantile at a time while setting others to default values, and then perform pairwise trade-offs between quantiles of different ranks to enhance efficiency. Specifically, we initialize the algorithm by setting $\beta_{r}=0$ for a selected rank $r$, with all other $\beta_{s}=1$ for $s\neq r$, reducing the non-conformity score vector to a scalar case. We incrementally increase $\beta_{r}$ by a small step size $\epsilon$ until the empirical coverage just exceeds the target level $1-\alpha$. Subsequently, we iteratively adjust $\beta_{r}$ and other $\beta_{s}$ by trading off small increments and decrements while maintaining the coverage constraint, accepting adjustments that improve efficiency. This process has a computational complexity of $\mathcal{O}(Kn)$ (or $\mathcal{O}(K^{2}n)$ when iterating over all positions) and empirically yields an approximate solution $\hat{\beta}$ that achieves superior efficiency compared to baseline methods. A demonstration of the coverage efficiency across iterations and the solution path of the approximated solution with different initializations is presented in Figure 2. The details of the algorithm can be found in Algorithm 2 in Appendix C.
Figure 2: Demonstration of the proposed approximated algorithm. Subplot 2(a) presents the efficiency of approximated solutions with different initializations toward optimal efficiency over iterations. Subplot 2(b) presents the solution paths of the approximated solutions for different initialization, with arrows indicating the path direction over iterations. The shaded area is the feasible region of the optimization problem, and the darker regions correspond to higher efficiency.
Corollary 3.3 (Validity of Approximated Solution).
For any approximated solution $\hat{\beta}$ output by Algorithm 2 and testing data $(X_{n+1},Y_{n+1})$ that is exchangeable with calibration $\mathcal{D}$, the prediction set $\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\hat{\beta})$ provides valid coverage, i.e.,
$\mathbb{P}(Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\hat{\beta}))\geq 1-\alpha.$
The proof of Theorem 3.3 is presented in Appendix B
3.3 Comparison with PCP
Probabilistic Conformal Prediction (PCP) (Wang et al.,, 2023) first studied CP for generative models and utilized the information of random samples. In contrast to our method, which considers all the distances between each random sample and the true response, PCP defines its non-conformity score as the minimum distance between random samples and the true response:
$E_{i}=\min_{k=1,\dots,K}\|Y_{i}-\hat{Y}_{i,k}\|.$ (7)
for each calibration data $(X_{i},Y_{i})$. For a new test data $(X_{n+1},Y_{n+1})$, the prediction set of PCP is constructed as the union of $K$ regions:
$\hat{\mathcal{C}}(X_{n+1})=\cup_{k=1}^{K}\hat{\mathcal{C}}_{k},$ (8)
where each region $\hat{\mathcal{C}}_{k}$ is defined by
Here $Q(\cdot;\alpha)$ is the $1-\alpha$ quantile of the empirical distribution of the non-conformity scores $E_{1},\dots,E_{n}$.
As noted, PCP adopts a scalar non-conformity score, which assigns a uniform radius to all prediction regions $\hat{C}_{k}$. This uniform treatment of random samples limits adaptability, as ideally, high-density regions should have larger radius, while low-density regions require smaller ones. In contrast, GCP-VCR introduces a vectorized non-conformity score that captures the full complexity of the prediction error distribution. By ranking samples based on their empirical density, our method enables individualized optimization of the ball radius for each sample. As demonstrated in Figure 4, the ball radius for low-density regions nearly diminishes, highlighting the flexibility of our approach. Another limitation of PCP arises when $K$ is small. In such cases, the empirical distribution of minimum distances may become heavy-tailed, requiring a large quantile (radius) to maintain coverage, leading to a conservative prediction set $\hat{\mathcal{C}}(X_{n+1})$. In contrast, GCP-VCR overcomes this by introducing an individualized quantile selection, relaxing the constraint for attaining the $1-\alpha$ quantile uniformly across all samples at each rank and reducing the impact of the distribution’s tail. As shown in Figure 2(b), where the target $\alpha=0.1$, but neither $\beta_{1}$ nor $\beta_{2}$ needs to be smaller than $\alpha$ to achieve the optimal efficiency.
@@ 4 Experiments
In this section, we evaluate the empirical performance of GCP-VCR compared with several state-of-the-art baselines on five synthetic data, one semi-synthetic data, and four real-world datasets. We demonstrate the superior efficiency and coverage validity of GCP-VCR, especially in handling multimodal and heterogeneous distributions.
@@ 4.1 Synthetic Data
We present the summary of the synthetic data results averaged over 100 experiments in Table 1 and highlight the most efficient method for each data. Overall, all methods achieve valid coverage rates, while GCP-VCR produces the most efficient prediction set. The visualizations of the prediction sets of different methods on these data are presented in Appendix E.1. We can observe that PCP and GCP-VCR can consistently provide discontinuous prediction sets and, thus are much more efficient than other methods for multimodal settings. Compared with PCP, GCP-VCR provides a more flexible and locally-adaptive prediction set, i.e., the coverage region of GCP-VCR is larger in the high-density region and smaller in the low-density region, especially in the Mix-Gaussian setting. This is the benefit that vectorized non-conformity scores provide compared with a scalar.
@@
4.2 MNIST: Digit Image Generation
We design a semi-synthetic experiment with the MNIST image dataset (Deng,, 2012) to assess the applicability and interpretability of GCP-VCR to high-dimensional data. Conformal prediction in high-dimensional data, such as images, texts, and audio, presents a unique challenge for interpreting prediction sets, as it is inapplicable to iterate over the entire space of target variables.
Consider the setting that given $X\in\{0,\dots,9\}$, the corresponding Y is an image of digit $X$. We train a variational-autoencoder (VAE) and create the prediction set in the latent embedding space. For the conditional sampling process, we assume access to an independent sampler that generates random images of digit $X$ not presented in the training/calibration/testing process. We then embed the sampled images with the trained encoder and construct the prediction set in the embedding space. Ahead of training, we will split part of the dataset for the sampler so that it is independent of the following process. The reason that we design this sampler instead of training a conditional VAE (cVAE) is that the embedding space of VAE tends to be more clustered than cVAE, thus the prediction set can be better demonstrated for illustrative purposes.
Figure 3 shows samples from the prediction set provided by PCP and GCP-VCR with $K=40$ conditioning on different $X$ values in each row. For each $X$, we randomly sample 9 images from the prediction set of PCP and GCP-VCR. Both of the methods achieve valid coverage rates. We can observe that the sampled images of GCP-VCR align closer with the target image for each $X$, indicating a tighter and more efficient prediction set of GCP-VCR. Quantitatively, as it is not applicable to directly measure the size of the prediction set, we propose to evaluate the miscoverage rate. It measures the empirical probability of an incorrect image falling into the prediction set. Under $K=40$, GCP-VCR attains an average miscoverage rate of $0.16\pm 0.02$ (standard error) over 30 experiments, compared to $0.22\pm 0.03$ for PCP. In Appendix E.2, we also present the comparison of the miscoverage rate of PCP and GCP-VCR under different choices of $K$. It shows that GCP-VCR consistently attains a lower miscoverage rate than PCP.
Figure 3: Comparison of MNIST images sampled from prediction sets of GCP-VCR and PCP, respectively. Each column corresponds to different values of $X$. The randomly sampled images of GCP-VCR align more closely with the target $X$ compared to PCP, indicating tighter prediction sets with fewer incorrect images.
(a) GCP-VCR
@@ 4.3 Real Data
We apply GCP-VCR along with baseline methods on four real-world datasets: blog feedback (Blog) Buza, (2014), facebook comment volumne Singh, (2015) variant 1 (Fb1) and variant 2 (Fb2), and New York City bike sharing (Bike). For each dataset, we randomly sample $30,000$ samples for training, $1,000$ samples for calibration and $1,000$ for testing. We set $K=50$ for the Bike dataset, and $K=20$ for the rest datasets. The summary of the real data results averaging over 30 experiments is presented in Table 2. Overall, GCP-VCR achieves the highest efficiency while maintaining the coverage rate. Notably, the efficiency gains of GCP-VCR are more pronounced on real datasets compared to synthetic data, which we attribute to such a difference in data complexity. Real datasets, such as the blog dataset with 280 features, exhibit highly complex conditional distributions $Y|X$, highlighting GCP-VCR’s efficiency and robustness in handling intricate data structures. In contrast, synthetic data with only a single feature, inherently has lower distributional complexity, leaving limited room for improvement. Additionally, we observe that for the first three datasets, R2CCP suffers from undercoverage issues and exhibits high variance in coverage rates across experiments.
We present a visual demonstration of the prediction set generated by our method, PCP, CHR and R2CCP in Figure 4, respectively. The blue region indicates the prediction set by each method and the red star marker indicates the target location. Both PCP and our method can deliver discontinuous coverage regions, whereas CHR and R2CCP provide a rectangle-shaped continuous set. It is also noted that, compared with PCP, GCP-VCR assigns larger circles’ radius to high-density regions while nearly vanishing the circles’ radius in low-density regions. Specifically, the circle at the top of the map lies in a low-density region, it almost shrinks to a point and disappears for GCP-VCR, whereas it remains large for PCP. Thus, GCP-VCR leads to significant improvements in flexibility and efficiency of prediction set.
@@ 5 Discussion and Conclusion
We introduced GCP-VCR, a novel conformal prediction method for generative models that leverages vectorized non-conformity scores with ranked samples to improve prediction set efficiency. By ranking conditional samples, our method captures nuanced error distribution and explores structural relationships among samples. Optimizing each entry of the non-conformity score vector separately enables a more flexible and data-adaptive prediction set. Specifically, GCP-VCR constructs discontinuous prediction sets as a union of regions centered at random samples, where the radius of each region is adapted to the model’s uncertainty level of each sample. We also proposed a computationally efficient algorithm to approximate the optimal quantile vector, ensuring scalability to large datasets.
Through extensive experiments on both synthetic and real-world datasets, GCP-VCR consistently demonstrated superior efficiency while maintaining coverage validity. In the MNIST experiment, we showcased the potential power of GCP-VCR in high-dimensional data, illustrating its ability to improve the reliability of predictions in modern high-stakes applications.
@@ Appendix A Average $m$-Nearest Neighbor Distance
Appendix B Proof
B.1 Proof of Theorem 3.1
@@ Appendix B Proof
We can rewrite the constraint in (6) as
$\begin{split}\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\beta)&=\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\max_{r=1,\dots,K}\mathbbm{1}\left\{E_{i,r}\leq Q_{r}(\beta_{r})\right\}\\ &=\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}(1-\mathbbm{1}\{E_{i}\succ Q(\beta)\})\\ &\geq 1-\alpha,\end{split}$ (10)
where for notation simplicity we let $Q(\beta)=(Q_{1}(\beta_{1}),\dots,Q_{K}(\beta_{K}))^{\top}$ denote the vector of quantiles and $E_{i}=(E_{i,1},\dots,E_{i,r})^{\top}$ for $i=\{1,\dots,n,\infty\}$.
For a set of vectors $v_{1},\dots,v_{n}\in\mathbb{R}^{p}$, we define the empirical multivariate quantile function as
$\hat{Q}(\{v_{1},\dots,v_{n}\};1-\alpha)$ $=\operatorname*{argmin}_{\beta\in\mathbb{R}^{K}}\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}$ (11)
s.t. $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{v_{i}\succ Q(\beta)\}\leq\alpha$ (12)
We then observe that the minimizer $\beta^{*}$ can be expressed with the defined empirical multivariate quantile function as
$\beta^{*}=\hat{Q}(\{E_{1},\dots,E_{n}\}\cup\{\infty\},1-\alpha).$
For a testing data $(X_{n+1},Y_{n+1})$ and its non-conformity score vector $E_{n+1}$, from Lemma 1 in Tibshirani et al., (2019) we have the following equivalence,
$E_{n+1}\succ\hat{Q}(\{E_{1},\dots,E_{n}\}\cup\{\infty\},1-\alpha)\Longleftrightarrow E_{n+1}\succ\hat{Q}(\{E_{1},\dots,E_{n},E_{n+1}\},1-\alpha).$ (13)
By the exchangeability assumption and the equivalence, we have $\mathbb{P}(E_{n+1}\succ Q(\beta^{*}))=\cfrac{\lfloor\alpha(n+1)\rfloor}{n+1}\leq\alpha.$ We observe that its complement set achieves the target coverage, i.e,
@@
$\cup_{r=1}^{K}\{E_{n+1,r}\leq Q_{r}(\beta^{*}_{r})\}\Longleftrightarrow Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*}).$
@@
B.2 Proof of Corollary 3.3
Proof.
We first observe that for any $\hat{\beta}$ output by algorithm 2, it satisfies that
$\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\hat{\beta})\geq 1-\alpha,$
which implies that $\sum\limits_{i=\{1,\dots,n,\infty\}}\mathbbm{1}\{E_{i}\succ Q(\hat{\beta})\}\leq\alpha$ by the derivation in equation (10). It implies that the rank of $Q(\hat{\beta})$ is of the top $\lfloor\alpha(n+1)\rfloor$ among $\{E_{1},\dots,E_{n},E_{\infty}\}$. By the exchangeability assumption and equation (13), $Q(\hat{\beta})$ is also of the top $\lfloor\alpha(n+1)\rfloor$ among $\{E_{1},\dots,E_{n},E_{n+1}\}$. We thus have $\mathbb{P}(E_{n+1}\succ Q(\hat{\beta}))=\frac{\lfloor\alpha(n+1)\rfloor}{n+1}\leq\alpha$ and its complement achieves the target coverage. The following steps are similar to the rest of the proof of Theorem 3.1.
Remark B.2.
Note that any $\beta$ satisfying the optimization constraint in (6) provides a valid coverage. In the multivariate case, there can exist a set of vectors that satisfy the constraint equation. Our proposed approximation algorithm searches within this set of valid vectors, ensuring that any output from the algorithm is valid. The distinction among these valid vectors lies in their corresponding efficiency, i.e., the objective function in (6). The minimizer $\beta^{*}$ achieves the highest efficiency, and the approximated solutions progressively approach this minimizer through iterations, as demonstrated in Figure 2.
@@ Appendix C Efficient Algorithm for Quantile Optimization
The key idea of the approximation method is motivated by the “trade-off” between quantiles for different ranks. Intuitively, if one quantile is decreased then another quantile has to be increased to maintain coverage validity. During this trade-off, the efficiency changes accordingly. We present a simple demonstration of this idea for $K=2$ in Figure 5. The colored region represents the feasible region in (6) with darker color indicating higher efficiency, and the red point shows the position of $\beta^{*}$. We can observe ($i$) a step-wise pattern on the boundary of the feasible region and ($ii$) $\beta^{*}$ appears on the boundary. The step-wise pattern corresponds to the trade-off between $\beta_{1}$ and $\beta_{2}$, which leads to different efficiency. Additionally, the feasible region is in a convex shape, also indicating the trade-off relationship between quantiles. Thus to approximate the optimal solution, we only need to find the boundary of the feasible region, and iteratively do the trade-offs between different quantiles.
Figure 5: Feasible region and optimal solution of optimization problem in (6) for $K=2$. The colored region shows the feasible region of $\beta$ with darker color indicating higher efficiency and the red point shows the position of $\beta^{*}$.
To be more specific, the algorithm initializes a $\beta_{k}=0$ and sets all the other entries of $\beta$ to be 1. Then, we gradually increase $\beta_{k}$ until the empirical coverage is just above the target coverage. The intuition behind such initialization is that it reduces the non-conformity score vector to the simple scalar case, as observed in the following equation,
@@
for $\beta_{k}\neq 1$ and $\beta_{j}=1$ for all $j\neq k$. Thus there exists such $\beta_{k}$ that $\beta$ can still satisfy empirical coverage with only $\beta_{k}\neq 1$. Let $\epsilon>0$ denote a small increment, we then randomly pick another entry $\beta_{j}$ for $j\neq k$ and do the trade-off between $\beta_{j}$ and $\beta_{k}$. We set $\beta_{k}=\beta_{k}+\epsilon$, and decrease $\beta_{j}$ until the empirical coverage is satisfied. Elementally, this process mimics the step-wise pattern, where we increase $\beta_{k}$ and decrease $\beta_{j}$ to do the trade-off and examine if the efficiency is improved. We iterate this process for different $j$ until reach the budget. We show that empirically, the approximated solution $\hat{\beta}$ still achieves superior efficiency than baseline methods. We refer to Algorithm 2 for the complete details of the empirical approximation method. As a remark for practice, we set $\epsilon=1/(n+1)$ as default. But in cases, if the quantiles are not updated after several iterations, one may consider increasing $\epsilon$ for a larger trade-off.
@@
3:  while $\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\beta)\geq 1-\alpha+\frac{1}{n+1}$ do
4:     $\beta_{k}\leftarrow\beta_{k}+\frac{1}{n+1}$;
@@
11:     while $\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\beta)<1-\alpha$ do
@@
19:        Set $\beta_{k}$ and $\beta_{j}$ as the $k$-th and $j$-th entry of $\beta$;